# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

### v2

@@ 4 Experiments
Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{\hat{d}}$. The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation. More specifically, the ViT first transforms inputs $x$ to tokens $t_{1},\ldots,t_{N}$ before further processing by the attention layers. The number of tokens $N$ is a constant defined by $(\frac{I}{P})^{2}$, where $I$ and $P$ are the image and patch sizes, respectively. Let ${\mathcal{T}}$ be the VWT associated with a vocabulary $v\in{\mathcal{V}}$, where $v$ is a visual word learned from some dataset $D$. The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$. In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency. We focus on the zero-shot setting by eschewing any form of end-to-end fine-tuning for $f$.
@@ 4.1 Datasets and Settings
We conduct our analysis through the lens of (i) classification performance of visual recognition, (ii) subgroup robustness, and (iii) generative performance of visual captioning. For (i) and (ii), we utilize three publicly available datasets (Waterbirds [63], CelebA [36], MetaShift [31]) that are typical benchmarks in robustness and fairness research [50, 35, 70]. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels4 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}. Further details on the defined subgroups are provided in Appendix 6.1. For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [29]. Following Huang et al. [25], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively. The former has 214 classes, while the latter has 200. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the encoded text label [^4] The prefix ”a photo of a ” is also added to encode each text label.. For (iii), we utilize the Karpathy test split of COCO dataset [33] and a validation set of NoCaps dataset [1] following the setting in previous work [30]. Finally, to study inference efficiency, we utilize all datasets for the visual tasks (i)-(iii) described above. We compute the efficiency per sample and average across all samples. For the wattage calculation, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs. The experiments are also conducted in an isolated environment, thus ensuring the efficiency measurements are as accurate as possible. Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations as this process is only done once and may be reused multiple times.
@@ 4.2 Implementation Details
For image classification, we load the pre-trained CLIP [46] model from HuggingFace [^5] https://huggingface.co/openai/clip-vit-base-patch16. An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP. For image captioning, we load the pre-trained BLIP [30] model from HuggingFace [^6] https://huggingface.co/Salesforce/blip-image-captioning-base. To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively. An image size of 384 $\times$ 384 is used with bicubic interpolation. Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise. Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [11] and token merging [5] as additional baselines. We denote the former as $Q_{8}$ and the latter as $ToME$. Following Bolya et al. [5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes. For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$. For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where ${\mathcal{V}}$ is the size of the vocabulary. Lastly, our experiments are conducted using a single NVIDIA A100 GPU. Since our focus is on the online setting (real-time), we set the batch size to 1 unless stated otherwise.
Figure 5: Visualization of long-form captioning on COCO. Longer captions are generated via a length penalty of 2.0 and a maximum length of 40. Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its lesser compression.
4.3 VWTs and Inference Efficiency
First we analyze the effects of VWTs on token length. We seek to understand how the choice of pre-processing data and vocabulary size affects the degree of compression. Table 1 shows the token length per sample (including [CLS]) on different datasets. Unlike ${\mathcal{T}}_{intra}^{0.5}$, the sequence lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal. First, we compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [10]. The in-domain dataset is represented by the training split if available. On text data, in-domain tokenizers [18, 19, 9, 67, 43] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution. Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K. Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K. Second, unlike text tokenizers, decreasing compression is seen as vocabulary size increases. With text, larger vocabularies ensure that more tokens are kept as words rather than subwords. We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
Having analyzed the effects on token length, we turn to the practical metrics of wattage and runtime. For online inference, savings in wattage should not incur a significant cost to runtime. In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToME$. Note that we compute both metrics using the image encoder of CLIP or BLIP only. First, we find the wattage of VWTs to be lower than $Base$ across the datasets. On COCO, this reduction is up to 25% with ${\mathcal{T}}_{inter}^{100}$. Naturally, efficiency decreases as vocabulary size increases due to smaller compression. Although $Q_{8}$ and $ToME$ result in a lower wattage than VWTs, we note the drawbacks in runtime. In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized. Meanwhile, $ToME$ can increase runtime by up to 100% on Waterbirds, CelebA, MetaShift, and OpenImages. We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{10000}$ on the aforementioned datasets. ${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{100}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ do so on COCO and NoCaps. We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
@@ 4.3 VWTs and Inference Efficiency
Table 4: Token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. Visual words are formed from patches after the pre-trained convolution layer of CLIP or BLIP. Unlike Table 1, poor compression is seen due to dissimilarity between patches and new visual words.
4.4 VWTs and Visual Performance
Another important factor in compression is the effect on model performance. In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively. For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by Sagawa et al. [50] and Romiti et al. [49]. First, we find the degradation in average accuracy to be small for the VWTs. The largest drop of up to 2% only is shown by ${\mathcal{T}}_{intra}^{0.5}$ on MetaShift. Likewise, both $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets. Second, we observe possible improvements in subgroup robustness with VWTs. On Waterbirds and CelebA, the worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively. Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToME$ relative to $Base$. Like Gee et al. [20], we find compression to not always be harmful to subgroup robustness by improving the WGA at a negligible cost to overall performance.
To further validate performance, we conduct additional evaluations on the large-scale OpenImages v6 dataset for zero-shot classification. We report mean Average Precision (mAP) as done by Huang et al. [25]. Likewise, from Tables 2 and 3, we find the VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting while retaining comparable performance to $Base$.
For image captioning with BLIP, we evaluate our models following the setting in Li et al. [30] by using the BLEU, CIDEr, and SPICE metrics w.r.t. the ground truth captions. On COCO, the VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$. On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{100}$. However, overall performance of ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ are still higher than $ToME$. Only $Q_{8}$ displays a slight improvement over $Base$. Like $ToME$, the VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
The generated captions by BLIP [30] can be used as priors for further training of other models. As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data. In Figure 5, we visualize the long-form captions on COCO. To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40. All other generation settings are kept the same. With longer captions, the generation may degenerate into unnecessary repetitions on certain samples. Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
@@ 4.4 VWTs and Visual Performance
Table 5: Token length (including [CLS]) and performance of the inter-image approach with random merging. The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$. Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{inter}^{100}$ except for CelebA. Performance does not change much for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ since barely any compression occurs.
4.5 Visualization of the Visual Words
To better understand the VWT, we visualize the patch matching of ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively. For the latter, we highlight in identical colors patches that are matched with one another. With ${\mathcal{T}}_{intra}^{0.5}$, the patches with the lowest pixel variance typically correspond to uninformative backgrounds. In most cases, the dropping of such patches continues to preserve the foreground object.
With ${\mathcal{T}}_{inter}^{100}$, we find that patches representing the background are more frequently grouped than those of the foreground object. On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background. On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals. We also observe that patch matching is capable of grouping similar but non-adjacent visual concepts. In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively. Our analysis shows that patch matching serves as a rudimentary form of image segmentation by identifying similar visual concepts for compression.
In Figure 4, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words. We show patches from ImageNet-1K (i.e. pre-processing data) that are matched to each visual word using the Euclidean distance. Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster. We observe that visual words depict basic features such as colors or edges. These features are also formed as an average representation of the patches that belong to each cluster. By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
Dataset Thresh Length Average $\uparrow$ Worst $\uparrow$
Waterbirds 0.2 97 76.71 24.09
0.3 79 74.85 21.13
0.4 66 73.35 16.30
0.5 58 72.81 13.14
CelebA 0.2 66 89.44 56.48
0.3 55 88.85 58.70
0.4 50 88.31 57.04
0.5 48 88.21 55.74
MetaShift 0.2 84 86.16 77.46
0.3 69 81.16 67.77
0.4 60 77.73 60.65
0.5 55 74.75 56.08
@@ 4.5 Visualization of the Visual Words
Table 6: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds. Higher thresholds can reduce sequences to 48 tokens on CelebA. Greater compression results in higher degradation in performance, particularly in subgroup robustness.
4.6 Ineffectiveness of the Embedding Space
In Table 4, we analyze the sequence compression using visual words formed in the embedding space. Instead of initializing the vocabulary by clustering patches in the pixel space, we do so with patches after the pre-trained convolution layer of CLIP or BLIP. During inference, we match the patches after the pre-trained convolution layer with these new visual words. Compared to Table 1, we observe a notable reduction in the compression irrespective of pre-processing data and vocabulary size. In Tang et al. [55], embeddings are shown to become progressively more similar to one another due to self-attention. As such, it is unsurprising that matching patches and visual words in the initial embedding space is found to be ineffective.
4.7 Random Merging of Tokens
In Table 5, we study the effects of randomly merging the tokens. We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$. First, we find the token length (including [CLS]) to differ noticeably from Table 1. For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets. Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression. Second, performance is shown to change from Table 3. We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift. For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs. With random merging, the captions are shown to deviate completely from those of $Base$ in Figure 6, thus further demonstrating that visual words are an effective criterion for grouping patches as shown in Figure 1.
4.8 Ablation of the Similarity Threshold
We have shown how the inter-image approach can improve the efficiency of online inference. To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 6. We seek to determine if exploiting higher thresholds for increased compression is a viable strategy. Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged. For Waterbirds and MetaShift, the WGA degrades more significantly than the average, especially with the former. Interestingly, average accuracy remains relatively unchanged while WGA improves significantly on CelebA irrespective of similarity threshold. We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.

### v4

@@ 4 Experiments
Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{\hat{d}}$. The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation. More specifically, the ViT first transforms inputs $x$ to tokens $t_{1},\ldots,t_{N}$ before further processing by the attention layers. The number of tokens $N$ is a constant defined by $(\frac{I}{P})^{2}$, where $I$ and $P$ are the image and patch sizes, respectively. Let $\mathcal{T}$} be the VWT associated with a vocabulary $v\in{\mathcal{V}}$, where $v$ is a visual word learned from some dataset $D$. The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$. In our experiments, we seek to analyze the effect of $\mathcal{T}$} for online inference. We focus on the zero-shot setting by eschewing any form of end-to-end fine-tuning for $f$.
Model Waterbirds CelebA MetaShift OpenImages
${\mathcal{T}}_{intra}^{0.5}+Q_{8}$ 76.31 33.65 89.43 52.41 93.14 87.78 65.29 59.27
${\mathcal{T}}_{inter}^{100}+Q_{8}$ 78.88 27.57 90.17 54.44 90.58 83.72 62.75 57.99
${\mathcal{T}}_{inter}^{1000}+Q_{8}$ 80.01 25.60 91.05 50.93 93.67 86.67 66.15 60.48
${\mathcal{T}}_{inter}^{10000}+Q_{8}$ 80.28 25.29 90.26 47.96 94.58 86.67 69.07 62.37
(a) CLIP (image classification and subgroup robustness)
Model COCO NoCaps
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
${\mathcal{T}}_{intra}^{0.5}+Q_{8}$ 32.45 103.18 101.98 14.46 94.93 13.74 87.13 12.78 94.36 13.66
${\mathcal{T}}_{inter}^{100}+Q_{8}$ 30.88 97.48 95.97 13.94 90.27 13.26 80.02 12.39 89.01 13.19
${\mathcal{T}}_{inter}^{1000}+Q_{8}$ 31.72 100.23 99.69 14.02 94.35 13.49 85.69 12.57 93.36 13.39
${\mathcal{T}}_{inter}^{10000}+Q_{8}$ 32.98 104.15 101.77 14.34 98.75 14.02 91.55 13.21 97.72 13.91
(b) BLIP (image captioning)
Table 4: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP). The VWTs are applied jointly with 8-bit quantization. Performance is shown to be similar to those with VWTs only, thereby displaying the mutual compatibility of VWTs with other compression techniques.
@@ 4.1 Datasets and Settings
We conduct our analysis through the lens of (i) classification performance of visual recognition, (ii) subgroup robustness, and (iii) generative performance of visual captioning. For (i) and (ii), we utilize three publicly available datasets (Waterbirds [wah2011caltech], CelebA [liu2015deep], MetaShift [liang2022metashift]) that are typical benchmarks in robustness and fairness research [sagawa2019distributionally, liu2021just, yang2023change]. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels5 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}. Further details on the defined subgroups are provided in Appendix 6.1. For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [kuznetsova2020open]. Following huang2023open, the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively. The former has 214 classes, while the latter has 200. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the encoded text label [^5] The prefix ”a photo of a ” is also added to encode each text label.. For (iii), we utilize the Karpathy test split of COCO dataset [lin2014microsoft] and a validation set of NoCaps dataset [agrawal2019nocaps] following the settings in previous work [li2022blip]. Finally, to study inference efficiency, we utilize all the datasets for the visual tasks (i)-(iii) described above in computing the energy (joule) consumed per sample. For calculating the power variable of energy, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs. The experiments are also conducted in an isolated environment, thus ensuring the efficiency measurements are as accurate as possible. Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations, as this process is only done once and may be reused multiple times.
@@ 4.2 Implementation Details
For image classification, we load the pre-trained CLIP [radford2021learning] model from HuggingFace [^6] https://huggingface.co/openai/clip-vit-base-patch16. An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP. For image captioning, we load the pre-trained BLIP [li2022blip] model from HuggingFace [^7] https://huggingface.co/Salesforce/blip-image-captioning-base. To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively. An image size of 384 $\times$ 384 is used with bicubic interpolation. Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise. Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [dettmers2022gpt3] and token merging [bolya2022token] as additional baselines. We denote the former as $Q_{8}$ and the latter as $ToMe$. Following bolya2022token, we utilize a reduction per layer for $ToMe$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes. For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$ as we found it to work best. Additional dropping ratios are explored in Table 10 of Appendix 7.1. For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where $\mathcal{V}$} is the size of the vocabulary. Lastly, our experiments are conducted using a single NVIDIA A100 GPU. Since our focus is on the online setting (real-time), we set the batch size to 1 unless stated otherwise.
@@
Table 5: Ablation of token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. Visual words for the inter-image approach are formed in the embedding space from patches after the pre-trained convolution layer of CLIP or BLIP. Unlike Table 1, poor compression is seen due to dissimilarity between the patches and new visual words during inference.
4.3 Experimental Results
VWTs and Inference Efficiency.
We begin by analyzing the effects of VWTs on token length to understand how the choice of pre-processing data and vocabulary size affects the degree of compression. Table 1 shows the token length per sample (including [CLS]) on different datasets. First, we observe a significant reduction in token length by the VWTs relative to $Base$. This reduction is most apparent for datasets with larger image sizes such as COCO and NoCaps. On the former, the token length is reduced by up to 54% with ${\mathcal{T}}_{inter}^{100}$. Second, the token lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal unlike ${\mathcal{T}}_{intra}^{0.5}$. We compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [deng2009imagenet]. The in-domain dataset is represented by the training split if available. On text data, in-domain tokenizers [gee2022fast, gee2023multi, dagan2024getting, yamaguchi2024empirical, minixhofer2024zero] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution. Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K. Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K. Third, unlike text tokenizers, decreasing compression is seen as vocabulary size increases. With text, larger vocabularies ensure that more tokens are kept as words rather than subwords. We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression. Note that henceforth, we use ImageNet-1K as our pre-processing data for the inter-image approach.
@@ 4.3 Experimental Results
Table 6: Ablation of token length (including [CLS]) and performance of the inter-image approach with random merging. The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$. Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{inter}^{100}$ except for CelebA. Performance does not change significantly for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ as negligible compression occurs.
Having analyzed the effects on token length, we turn to the practical metrics of energy. In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToMe$. Note that we compute both metrics using the image encoder of CLIP or BLIP only. First, we find the energy consumed by VWTs to be generally lower than $Base$ across the datasets. On OpenImages (Common) and COCO, this reduction is up to 43% with ${\mathcal{T}}_{intra}^{0.5}$ and 47% with ${\mathcal{T}}_{inter}^{100}$, respectively. Interestingly, the inter-image approach appears to be most effective for datasets with larger image sizes (384 $\times$ 384) as shown with COCO and NoCaps rather than the remaining datasets with smaller image sizes (224 $\times$ 224). This efficiency also decreases as the vocabulary size increases due to smaller compression. For $Q_{8}$ and $ToMe$, we observe significant drawbacks in energy efficiency relative to $Base$. In the case of the former, the increase in energy consumed can reach as high as 500% or more on Waterbirds, CelebA, and MetaShift. This can be explained by the significantly longer runtime of $Q_{8}$ as tensors need to be repeatedly quantized and dequantized. Meanwhile, $ToMe$ can increase energy consumption by up to 50% or more. We have shown how VWTs are more suitable for online inference by improving energy efficiency, particularly for datasets with larger image sizes. As mentioned previously, we do not consider the Pre-Process step of the inter-image approach in our efficiency calculations as it is only executed once and not repeated during inference. This is similar to how model efficiency for large language models is calculated by excluding the construction of the tokenizer’s vocabulary via Byte-Pair Encoding. For reference, this step on the CPU takes approximately 30 minutes to an hour, depending on the vocabulary size [^8] Note that this pre-processing is done on ImageNet-1K. Naturally, dataset size also affects the total duration..
VWTs and Visual Performance.
Another important factor in compression is the effect on model performance. In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively. For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by sagawa2019distributionally and romiti2022realpatch. To further validate performance, we report the mean Average Precision (mAP) on the large-scale OpenImages v6 dataset [huang2023open]. For image captioning with BLIP, we evaluate our models following the setting in li2022blip by using the BLEU, CIDEr, and SPICE metrics w.r.t. the ground truth captions.
First, we find the degradation in performance relative to $Base$ to be smallest for $Q_{8}$ across the datasets. However, this comes at the cost of a significant rise in energy consumed as shown in Table 2. We also observe that VWTs are competitive with $ToMe$, particularly on the image captioning tasks of COCO and NoCaps. Second, both $Q_{8}$ and the VWTs are shown to improve the subgroup robustness on Waterbirds and CelebA. The worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively. Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToMe$. Like gee2023compressed, we find compression to not always be harmful to subgroup robustness by improving the WGA at a small or negligible cost to overall performance. Third, the intraintra-image approach is seen to work best on MetaShift and OpenImages while the interinter-image approach is more suitable on the remaining datasets. The average accuracy with ${\mathcal{T}}_{inter}^{100}$ drops by up to 5% and 7% on MetaShift and OpenImages, respectively. We posit that the lower performance may stem from the lack of uniform backgrounds (e.g., skies in Waterbirds) that causes compression to increasingly shift to the foreground object instead as shown in Figure 1.
Model Waterbirds CelebA MetaShift
Length Average $\uparrow$ Worst $\uparrow$ Length Average $\uparrow$ Worst $\uparrow$ Length Average $\uparrow$ Worst $\uparrow$
0.2 97 76.71 24.09 66 89.44 56.48 84 86.16 77.46
0.3 79 74.85 21.13 55 88.85 58.70 69 81.16 67.77
0.4 66 73.35 16.30 50 88.31 57.04 60 77.73 60.65
0.5 58 72.81 13.14 48 88.21 55.74 55 74.75 56.08
@@
Table 7: Ablation of token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds. Higher thresholds for the inter-image approach can reduce sequences to 48 tokens on CelebA. Greater compression results in higher degradation in performance, particularly in subgroup robustness.
Visualizing the VWTs.
To better understand the VWTs, we visualize the patch dropping of ${\mathcal{T}}_{intra}^{0.5}$ and patch matching of ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively. For the latter, we highlight in identical colors the patches that are matched with one another. With ${\mathcal{T}}_{intra}^{0.5}$, patches with the lowest pixel variance typically correspond to uninformative backgrounds. In most cases, dropping such patches continues to preserve the foreground object. With ${\mathcal{T}}_{inter}^{100}$, we find that patches representing the background are more frequently grouped than those of the foreground object. On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating spurious correlations with the background. On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths. We also observe that ${\mathcal{T}}_{inter}^{100}$ may group similar but non-adjacent visual concepts. In certain examples of MetaShift and NoCaps, multiple cats and foods are matched together, respectively. Our analysis shows that patch matching serves as a rudimentary form of image segmentation by identifying similar visual concepts for compression.
In Figure 4, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words. We show patches from ImageNet-1K (pre-processing data) that are matched to each visual word using the Euclidean distance. Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster. We find that visual words depict basic features such as colors or edges. These features are also formed as an average representation of the patches that belong to each cluster. By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
The generated captions by BLIP [li2022blip] can be used as priors for further training of other models. As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data. In Figure 5, we visualize the long-form captions on COCO. To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40. All other generation settings are kept the same. With longer captions, the generation may degenerate into unnecessary repetitions on certain samples. Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
4.4 Ablation Studies
VWTs with Quantization.
We designed the VWT to be a training-free approach to efficient online inference. As such, we exclude any form of fine-tuning [lin2024not, hao2024meft, simoulin2024memory] (e.g., knowledge distillation [tian2019contrastive, han2024amd, son2021densely]) for performance recovery. However, should it be desired, our method may be used alongside network compression [tian2019contrastive, han2024amd, son2021densely] (i.e., minimizing the Kullback–Leibler divergence between the softened outputs of the larger teacher and smaller student), memory-efficient fine-tuning [hao2024meft, simoulin2024memory], and other compression techniques (e.g., quantization) as the VWT only influences the initial input sequence to the vision transformer. In Table 4, we provide additional scores when VWTs are combined with 8-bit quantization. We find the performance to be similar to those in Table 3, thereby showing that VWTs are highly compatible with other compression approaches as mentioned above.